# Rotatable Antenna Enabled Wireless Communication: Modeling and Optimization

### v4

@@ Abstract
In this paper, we propose a new rotatable antenna (RA) model to improve the performance of wireless communication systems. Different from conventional fixed antennas, the proposed RA system can flexibly and independently alter the three-dimensional (3D) boresight direction of each antenna to achieve a desired array directional gain pattern. Specifically, we investigate an RA-enabled uplink communication system, where the receive beamforming and the boresight directions of all RAs at the base station (BS) are jointly optimized to maximize the minimum signal-to-interference-plus-noise ratio (SINR) among all the users. In the special single-user and free-space propagation setup, the optimal boresight directions of RAs are derived in closed form with the maximum-ratio combining (MRC) beamformer applied at the BS. Moreover, we analyze the asymptotic performance with an infinite number of antennas based on this solution, which theoretically proves that the RA system can achieve a higher array gain than the fixed-antenna system. In the general multi-user and multipath channel setup, we first propose an alternating optimization (AO) algorithm to alternately optimize the receive beamforming and the boresight directions of RAs in an iterative manner. Then, a two-stage algorithm that solves the formulated problem without the need for iteration is proposed to further reduce computational complexity. Simulation results are provided to validate our analytical results and demonstrate that the proposed RA system can significantly improve the communication performance as compared to other benchmark schemes.
Index Terms:
Rotatable antenna (RA), near-field modeling, array directional gain pattern, performance analysis, pointing vector optimization, antenna boresight, antenna orientation.
@@ I Introduction
In the rapidly evolving landscape of global information and communications technology (ICT), the forthcoming sixth-generation (6G) wireless network is envisioned to support even more densely connected users and devices. It will serve a wider range of applications and services, thus demanding significantly higher performance requirements compared to its preceding generations [1]. Undoubtedly, multiple-input multiple-output (MIMO) is one of the most critical technologies for the current fifth-generation (5G) mobile communication. MIMO can dramatically enhance the transmission rate and reliability of wireless networks. This is achieved through beamforming and multiplexing using multiple antennas at the transceivers [2]. However, the channel capacity and spectrum efficiency achieved by conventional MIMO are insufficient to meet the stringent requirements of 6G in its new applications. To further improve spatial resolution and degrees-of-freedom (DoFs), wireless networks tend to integrate drastically more antennas into arrays at the base stations (BSs), thereby evolving MIMO into massive MIMO, and ultimately into extremely large-scale MIMO [3, 4, 5, 6, 7].
@@
Although larger-scale MIMO can offer substantial array and spatial multiplexing gains, it comes at the expense of much higher hardware costs and power consumption. Furthermore, simply increasing the number of antennas cannot fully exploit the spatial DoFs, as traditional fixed antennas lack the flexibility to adjust their positions or orientations. Recently, fluid antenna system (FAS) and movable antenna (MA) have been proposed as promising technologies to overcome this limitation and have attracted growing attention in wireless communication [8, 9, 10]. Compared to the fixed-antenna architecture, FAS/MA enables the local movement of antennas in a specified region through different antenna movement mechanisms, which can proactively reshape the wireless channels to more favorable conditions and thus achieve higher capacity without increasing the number of antennas. Furthermore, with such a new DoF offered at the physical layer, it has been validated that FAS/MA can achieve various significant performance advantages, including interference mitigation, flexible beamforming, and multiplexing enhancement [11, 12, 13]. By leveraging these capabilities of FAS/MA, substantial efforts have been devoted to integrating them with cutting-edge wireless technologies, such as integrated sensing and communications (ISAC) [14], unmanned aerial vehicle (UAV) communications [15], intelligent reflecting surface (IRS) [16, 17, 18], and over-the-air computation [19]. Nevertheless, while FAS/MA can bring numerous performance advantages, their practical implementation is highly constrained by the response time and/or movement speed of the antennas. Additionally, existing works on FAS/MA still face limitations in terms of spatial flexibility and performance enhancement since only the positions of antennas are adjusted while their orientations are fixed. To fully exploit all six-dimensional (6D) spatial DoFs, 6D movable antenna (6DMA) has been recently proposed to flexibly adjust both the three-dimensional (3D) position and 3D rotation of distributed antennas/arrays [20, 21]. Based on the long-term/statistical user channel distribution, the 6DMA-equipped transceiver can adaptively allocate its antenna resources to improve the array and spatial multiplexing gains.
@@
Motivated by the above, we propose in this paper a new antenna architecture, called rotatable antenna (RA), as a simplified implementation of 6DMA to improve the performance of wireless communication cost-effectively. In the RA system, the 3D orientation/boresight of each directional antenna can be independently adjusted while its 3D position remains fixed. As such, unlike the translation movement in MA and 6DMA, which requires additional space such as sliding tracks and movable area, RA only requires local rotational adjustment, which can be more readily achieved through compact mechanical or integrated electronic design [22, 23]. This offers RA great scalability and compatibility with existing wireless systems. Compared to the conventional fixed-antenna architecture, RA can enhance communication and sensing performance substantially by flexibly adjusting the antenna orientation/boresight within the full 3D space. In this way, RA provides a practical solution for enhancing array gains in desired directions to boost the transmit/receive signal power, while reducing the radiation power in undesired directions to avoid information leakage and interference. Therefore, by strategically designing the beamforming and antenna orientation/boresight, the RA system can be deployed to further improve the array/multiplexing gain and enhance the sensing resolution/range in various applications, such as ISAC, massive machine-type communication (mMTC), simultaneous localization and mapping (SLAM), and others.
@@
  • Adjusting the antenna orientation/boresight requires judicious consideration of both the propagation environment and the antenna’s directional gain pattern. To facilitate this, we define a pointing vector to characterize the 3D orientation/boresight of each RA, and then construct a new multipath geometric near-field channel model based on the pointing vectors of all RAs. Under this channel model, we further formulate a minimum signal-to-interference-plus-noise ratio (SINR) maximization problem to jointly optimize the receive beamforming and the pointing vectors of all RAs.
@@
  • For the special single-user and free-space propagation setup with the maximum-ratio combining (MRC) beamformer applied at the BS, we derive the optimal pointing vectors of RAs in closed form. Meanwhile, we also derive a closed-form expression and lower/upper bounds for the signal-to-noise ratio (SNR) under the uniform linear/planar array (ULA/UPA) setting, which show that the resultant SNR first increases linearly with the number of antennas and eventually converges to a certain limit. Additionally, we present an asymptotic analysis as the number of antennas goes to infinity, theoretically demonstrating that a wider rotational range for antenna boresight adjustment enables the proposed RA system to exploit more spatial DoFs, thereby achieving a higher array gain.
@@
  • For the general multi-user and multipath channel setup, we formulate a minimum-SINR maximization problem to balance the array directional gains among users across their multipath channels. To address this problem, we first propose an alternating optimization (AO) algorithm that alternately optimizes the receive beamforming and the pointing vectors of RAs in an iterative manner until convergence is achieved. In particular, with the optimal minimum mean-square error (MMSE) beamformer applied at the BS, the subproblem that optimizes the pointing vectors of RAs is solved by the successive convex approximation (SCA) technique. To reduce computational complexity, we further propose a two-stage algorithm that solves a weighted channel power gain maximization problem based on the zero-forcing (ZF) beamformer without the need for iteration.
@@
  • Simulation results validate our theoretical analysis and demonstrate that the proposed RA system can significantly improve communication performance over various benchmark schemes. It is shown that even with a small rotational range for antenna boresight adjustment, the RA system can reap considerable performance gains over the fixed-antenna system. Furthermore, the performance advantages of RA in flexibly balancing the directional gain over multipath channels become more pronounced with stronger antenna directivity.
@@
The remainder of this paper is organized as follows. Section II introduces the system model and problem formulation for designing the RA-enabled wireless communication system. In Section III, we derive the optimal closed-form solution and analyze the asymptotic performance under the single-user setup. Section IV proposes the AO algorithm and the two-stage algorithm to solve the formulated problem under the multi-user setup. Section V presents simulation results to evaluate the performance of the proposed system and algorithms. Finally, we conclude the paper in Section VI.
@@
Notation: Upper-case and lower-case boldface letters denote matrices and column vectors, respectively. Superscripts $(\cdot)^{T}$, $(\cdot)^{H}$, and $(\cdot)^{-1}$ stand for the transpose, Hermitian transpose, and matrix inversion operations, respectively. The sets of $a\times b$ dimensional complex and real matrices are denoted by $\mathbb{C}^{a\times b}$ and $\mathbb{R}^{a\times b}$, respectively. $\mathcal{O}(\cdot)$ denotes the standard big-O notation. For a vector $\bm{x}$, $\|\bm{x}\|$ denotes its $\ell_{2}$-norm, $\mathrm{Re}\{\bm{x}\}$ denotes its real part, $\mathrm{diag}(\bm{x})$ returns a diagonal matrix with the elements in $\bm{x}$ on its main diagonal, and $[\bm{x}]_{a:b}$ denotes the subvector of $\bm{x}$ consisting of the elements from $a$ to $b$. For a matrix $\mathbf{X}$, $\mathrm{Tr}(\mathbf{X})$ and $\mathrm{rank}(\mathbf{X})$ denote its trace and rank, $[\mathbf{X}]_{a,b}$ denotes the $(a,b)$-th entry of matrix $\mathbf{X}$, $[\mathbf{X}]_{a:b,c:d}$ denotes the submatrix of $\mathbf{X}$ consisting of the elements located in rows $a$ to $b$ and columns $c$ to $d$, and $\mathbf{X}\succeq 0$ implies that $\mathbf{X}$ is positive semi-definite. $\mathbf{I}$ and $\mathbf{0}$ denote an identity matrix and an all-zero matrix, respectively, with appropriate dimensions. The distribution of a circularly symmetric complex Gaussian (CSCG) random vector with zero mean and covariance matrix $\bm{\Sigma}$ is denoted by $\mathcal{N}_{c}(\bm{0},\bm{\Sigma})$; and $\sim$ stands for “distributed as”.
@@ II System Model and Problem Formulation
where $\delta_{k}\triangleq\frac{\Delta}{r_{k}}$ and $n=1,2,\dots,N$. Note that $\delta_{k}\ll 1$ since the RA separation $\Delta$ is typically on the order of wavelength in practice.
@@ II-A Antenna Boresight Rotation
The original orientations/boresights of all RAs are assumed to be parallel to the $z$-axis, and the boresight direction of each RA can be independently adjusted in 3D space by mechanical or electronic methods [22]. Specifically, mechanical control typically utilizes servo motors or micro-electromechanical systems (MEMS) to physically rotate the orientations of directional antennas, thereby adjusting their boresight directions [24]. In contrast, electronic control retains fixed antenna orientations while enabling boresight rotation through electronic techniques such as multi-feed switching or PIN diode biasing [25]. It is noted that mechanical control generally offers a wider range for boresight rotation, whereas electronic control provides much faster response and better compatibility with existing wireless systems.
@@
where $f_{x,n}$, $f_{y,n}$, and $f_{z,n}$ are the projections of RA $n$’s pointing vector on the $x$-, $y$-, and $z$-axes, respectively. For the boresight rotation of RA $n$, we let $\theta_{\mathrm{z},n}$ represent its zenith angle (i.e., the angle between the boresight direction of RA $n$ and the $z$-axis) and $\theta_{\mathrm{a},n}$ represent its azimuth angle (i.e., the angle between the projection of the boresight direction of RA $n$ onto the $x$-$y$ plane and the $x$-axis). Accordingly, we have $f_{x,n}=\operatorname{sin}\theta_{\mathrm{z},n}\operatorname{cos}\theta_{\mathrm{a},n}$, $f_{y,n}=\operatorname{sin}\theta_{\mathrm{z},n}\operatorname{sin}\theta_{\mathrm{a},n}$ and $f_{z,n}=\operatorname{cos}\theta_{\mathrm{z},n}$. Furthermore, we have $\|\vec{\mathbf{f}}_{n}\|=1$ due to normalization. To account for practical rotational constraint and mitigate antenna coupling between any two RAs, the zenith angle of each RA should be confined to a specific range [26]:
@@
where $\theta_{\mathrm{max}}\in[0,\frac{\pi}{2}]$ is the maximum zenith angle that each RA is allowed to adjust. [^1] To characterize the fundamental performance of the proposed RA system, we assume that the orientation/boresight of each RA can be continuously tuned in its rotational range, subject to the constraint in (4), while in practice it is usually chosen from a finite number of available directions for the ease of hardware implementation. The design of RA systems with discrete orientation/boresight rotation will be left for future work.
@@ II-B Channel Model
We consider the scattering environment with $Q$ distributed scatterer clusters, where the position of scatterer cluster $q$ is represented by $\mathbf{c}_{q}\in\mathbb{R}^{3\times 1}$ with $q=1,2,\dots,Q$. In addition, we consider the narrow-band frequency-flat channel model for ease of exposition. [^2] The current narrow-band model can be extended to wideband scenarios [28, 29] by employing orthogonal frequency division multiplexing (OFDM) technology, which divides the wideband channel into multiple orthogonal sub-bands, each undergoing frequency-flat fading. For this extension, the beamforming and orientations/boresights of RAs need to be jointly optimized across all sub-bands to maximize their achievable sum rate. Based on the Friis Transmission Equation and the directional gain pattern adopted in (5), the channel power gain between user $k$ and RA $n$ can be modeled as [30]
$g_{\mathrm{u},k}(\vec{\mathbf{f}}_{n})$ $\approx\int_{\mathbb{A}_{n}}\frac{1}{4\pi\|\mathbf{u}_{k}-\mathbf{a}\|^{2}}G_{0}\left(\frac{\vec{\mathbf{f}}_{n}^{T}(\mathbf{u}_{k}-\mathbf{a})}{\|\mathbf{u}_{k}-\mathbf{a}\|}\right)^{2p}\mathrm{d}\mathbf{a}$
$=\frac{A}{4\pi r_{k,n}^{2}}G_{0}\operatorname{cos}^{2p}(\epsilon_{k,n}),$ (6)
where the integral space $\mathbb{A}_{n}=\left[n_{x}\Delta-\frac{\sqrt{A}}{2},n_{x}\Delta+\frac{\sqrt{A}}{2}\right]\times\left[n_{y}\Delta-\frac{\sqrt{A}}{2},n_{y}\Delta+\frac{\sqrt{A}}{2}\right]$ corresponds to the surface region of RA $n$, $\mathbf{a}\in\mathbb{A}_{n}$ represents any point on plane $\mathbb{A}_{n}$, and $\operatorname{cos}(\epsilon_{k,n})\triangleq\vec{\mathbf{f}}_{n}^{T}\vec{\mathbf{u}}_{k,n}$ is the projection between $\vec{\mathbf{f}}_{n}$ and $\vec{\mathbf{u}}_{k,n}$ with $\vec{\mathbf{u}}_{k,n}\triangleq\frac{\mathbf{u}_{k}-\mathbf{w}_{n}}{\|\mathbf{u}_{k}-\mathbf{w}_{n}\|}$ being the direction vector from RA $n$ to user $k$. Similarly, the channel power gain between scatterer cluster $q$ and RA $n$ is modeled as
$g_{\mathrm{c},q}(\vec{\mathbf{f}}_{n})\approx\frac{A}{4\pi d_{q,n}^{2}}G_{0}\operatorname{cos}^{2p}(\tilde{\epsilon}_{q,n}),$ (7)
where $d_{q,n}=\|\mathbf{c}_{q}-\mathbf{w}_{n}\|$ is the distance between scatterer cluster $q$ and RA $n$, and $\operatorname{cos}(\tilde{\epsilon}_{q,n})\triangleq\vec{\mathbf{f}}_{n}^{T}\vec{\mathbf{c}}_{q,n}$ is the projection between $\vec{\mathbf{f}}_{n}$ and $\vec{\mathbf{c}}_{q,n}$ with $\vec{\mathbf{c}}_{q,n}\triangleq\frac{\mathbf{c}_{q}-\mathbf{w}_{n}}{\|\mathbf{c}_{q}-\mathbf{w}_{n}\|}$ denoting the direction vector from RA $n$ to scatterer cluster $q$. Note that the channel power gains modeled in (6) and (7) account for both path gain and directional gain.
@@
$h_{k}^{\mathrm{NLoS}}(\vec{\mathbf{f}}_{n})$ $=\sum_{q=1}^{Q}{\frac{\sqrt{\sigma_{q}g_{\mathrm{c},q}(\vec{\mathbf{f}}_{n})}}{t_{k,q}}}e^{-j\frac{2\pi}{\lambda}(d_{q,n}+t_{k,q})+j\chi_{q}},$ (9)
where $\lambda$ is the signal wavelength, $\sigma_{q}$ represents the radar cross section (RCS) of scatterer cluster $q$, $\chi_{q}$ represents the phase shift introduced by scatterer cluster $q$, and $t_{k,q}=\|\mathbf{u}_{k}-\mathbf{c}_{q}\|$ denotes the distance between user $k$ and scatterer cluster $q$. Thus, by superimposing the LoS and NLoS channel components, the overall multipath channel between user $k$ and the BS is given by

### v5

@@ Abstract
In this paper, we propose a new rotatable antenna (RA) model to improve the performance of wireless communication systems. Different from conventional fixed antennas, the proposed RA system can flexibly and independently alter the boresight direction of each antenna via mechanical or electronic means to exploit new spatial degrees-of-freedom (DoFs). Specifically, we investigate an RA-enabled uplink communication system, where the receive beamforming and the boresight directions of all RAs at the base station (BS) are jointly optimized to maximize the minimum signal-to-interference-plus-noise ratio (SINR) among all the users. In the special single-user and free-space propagation setup, the optimal boresight directions of RAs are derived in closed form with the maximum-ratio combining (MRC) beamformer applied at the BS. In the general multi-user and multipath channel setup, we first propose an alternating optimization (AO) algorithm to alternately optimize the receive beamforming and the boresight directions of RAs in an iterative manner. Then, a two-stage algorithm that solves the formulated problem without the need for iteration is proposed to further reduce computational complexity. Moreover, we extend the channel model to incorporate polarization effects and frequency-selective fading while catering to antenna boresight rotation. Simulation results are provided to validate our analytical results and demonstrate that the proposed RA system can significantly improve the communication performance as compared to other benchmark schemes.
@@ I Introduction
In the rapidly evolving landscape of global information and communications technology (ICT), the forthcoming sixth-generation (6G) wireless network is envisioned to support extremely high user/device density and a wider range of applications and services. These developments impose significantly higher performance requirements than those of previous generations [34]. Undoubtedly, multiple-input multiple-output (MIMO) is one of the most critical technologies to dramatically enhance the transmission rate and reliability of wireless networks. However, the channel capacity and spectrum efficiency achieved by conventional MIMO are insufficient to meet the stringent requirements of 6G in its new applications. To further improve spatial resolution and degrees-of-freedom (DoFs), wireless networks tend to integrate drastically more antennas into arrays at the base stations (BSs), thereby evolving MIMO into massive MIMO, and ultimately into extremely large-scale MIMO [16, 46, 12, 14, 4].
@@
Although larger-scale MIMO can offer substantial array and spatial multiplexing gains, it comes at the expense of much higher hardware costs and power consumption. Furthermore, simply increasing the number of antennas cannot fully exploit the spatial DoFs, as traditional fixed antennas lack the flexibility to adjust their positions or orientations. Recently, fluid antenna system (FAS) and movable antenna (MA) have been proposed as promising technologies to overcome this limitation and have attracted growing attention in wireless communication [19, 36, 49]. Compared to the fixed-antenna architecture, FAS/MA enables the local movement of antennas in a specified region through different antenna movement mechanisms. This flexibility enables proactive reshaping of the wireless channels into more favorable conditions, thereby achieving higher capacity without increasing the number of antennas. Furthermore, leveraging the additional DoFs available at the physical layer, FAS/MA has been demonstrated to provide notable performance benefits in terms of interference mitigation, flexible beamforming, and multiplexing enhancement [50, 37, 20].
Building upon the aforementioned advantages, substantial efforts have been devoted to integrating FAS/MA with cutting-edge wireless technologies, such as integrated sensing and communications (ISAC) [22], intelligent reflecting surface (IRS) [25, 35, 45], and over-the-air computation [41]. Nevertheless, although FAS/MA can bring numerous performance advantages, their practical implementation is highly constrained by the response time and/or movement speed of the antennas. Additionally, existing works on FAS/MA still face limitations in terms of spatial flexibility and performance enhancement since only the positions of antennas are adjusted while their orientations are fixed. To fully exploit all six-dimensional (6D) spatial DoFs, 6D movable antenna (6DMA) has been recently proposed to flexibly adjust both the three-dimensional (3D) position and 3D rotation of distributed antennas/arrays [26, 29, 28, 27]. Based on the long-term/statistical user channel distribution, the 6DMA-equipped transceiver can allocate its antenna resources to improve the multi-user communication throughput.
@@
Motivated by the above, we propose in this paper a new antenna architecture, called rotatable antenna (RA), as a simplified realization of 6DMA to enhance the performance of wireless systems cost-effectively. In the RA system, the 3D orientation/boresight of each directional antenna can be independently adjusted via mechanical or electronic means while its 3D position remains fixed. As shown in Table I, unlike MA and 6DMA that necessitate additional space for translational movement, RA only involves local rotational adjustment, which can be readily achieved through compact mechanical or integrated electronic design [43, 44, 39]. This thus offers RA great scalability and compatibility with existing wireless systems. Compared to the fixed-antenna architecture, RA enhances communication and sensing performance by flexibly adjusting the antenna orientation/boresight. In this way, RA offers a practical and promising solution to enhance array gains towards desired directions while suppressing radiation towards undesired ones to reduce leakage and interference. Therefore, by strategically designing the beamforming and antenna orientation/boresight, the RA system can be deployed to further improve the array/multiplexing gain and enhance the sensing resolution/range in various applications.
TABLE I: Comparison of Different Flexible Antenna Architectures [43]
Antenna
Architecture
Hardware Implementation
Reconfigurable Parameter
Deployment
Complexity
System
Overhead
RA [43]
Mechanical or electronic means
Antenna 3D boresight direction
Very low Low
FAS [36] Pixel antennas or liquid materials
Antenna shape and 3D position
Moderate Moderate
MA [49] Motors Antenna 3D position Moderate Moderate
6DMA [27] Motors and flexible cables
Antenna 3D position and 3D rotation
Moderate to high Moderate to high
@@
  • We first introduce a pointing vector to characterize the 3D orientation/boresight direction of each RA, and incorporate the antenna boresight rotation into the multipath geometric-based channel model. The channel model is then further extended to account for polarization effects and frequency-selective fading while catering to antenna boresight rotation.
@@
  • For the special single-user and free-space propagation setup with the maximum-ratio combining (MRC) beamformer applied at the BS, we derive the optimal pointing vectors of RAs in closed form. Meanwhile, we also derive a closed-form expression and lower/upper bounds for the signal-to-noise ratio (SNR) under the uniform linear/planar array (ULA/UPA) setting, which show that the resultant SNR first increases linearly with the number of antennas and eventually converges to a certain limit.
@@
  • For the general multi-user and multipath channel setup, we formulate a minimum signal-to-interference-plus-noise ratio (SINR) maximization problem to balance the array directional gains among users. To address this problem, we first propose an alternating optimization (AO) algorithm that alternately optimizes the receive beamforming and the pointing vectors of RAs in an iterative manner until convergence is achieved. To reduce computational complexity, we further propose a two-stage algorithm that solves a weighted channel power gain maximization problem based on the zero-forcing (ZF) beamformer without the need for iteration.
@@
  • Simulation results validate our theoretical analysis and demonstrate that the proposed RA system can significantly improve communication performance over various benchmark schemes. It is shown that even with a small rotational range for antenna boresight adjustment, the RA system can achieve considerable performance gains over the fixed-antenna system. Furthermore, the performance advantages of RA in flexibly balancing the directional gain over multipath channels become more pronounced with stronger antenna directivity.
@@
The remainder of this paper is organized as follows. Section II introduces the system model and problem formulation for designing the RA-enabled wireless communication system. In Section III, we derive the optimal closed-form solution and analyze the asymptotic performance under the single-user setup. In Section IV, we propose the AO algorithm and the two-stage algorithm to solve the formulated problem under the multi-user setup. Simulation results are presented in Section V to evaluate the performance of the proposed system and algorithms. Finally, we conclude the paper in Section VI.
@@
Notation: Upper-case and lower-case boldface letters denote matrices and column vectors, respectively. Superscripts $(\cdot)^{T}$, $(\cdot)^{\ast}$, $(\cdot)^{H}$, and $(\cdot)^{-1}$ stand for the transpose, conjugate, Hermitian transpose, and matrix inversion operations, respectively. The sets of $a\times b$ dimensional complex and real matrices are denoted by $\mathbb{C}^{a\times b}$ and $\mathbb{R}^{a\times b}$, respectively. $\left\lfloor\cdot\right\rfloor$ is the floor function. $\mathcal{O}(\cdot)$ denotes the standard big-O notation. For a vector $\bm{x}$, $\|\bm{x}\|$ denotes its $\ell_{2}$-norm, $\mathrm{Re}\{\bm{x}\}$ denotes its real part, $\mathrm{diag}(\bm{x})$ returns a diagonal matrix with the elements in $\bm{x}$ on its main diagonal, and $\bm{x}]_{a:b}$ denotes the subvector of $\bm{x}$ consisting of the elements from $a$ to $b$. For a matrix $\mathbf{X}$, $\mathrm{Tr}(\mathbf{X})$ and $\mathrm{rank}(\mathbf{X})$ denote its trace and rank, $\mathbf{X}]_{a,b}$ denotes the $(a,b)$-th entry of matrix $\mathbf{X}$, $\mathbf{X}]_{a:b,c:d}$ denotes the submatrix of $\mathbf{X}$ consisting of the elements located in rows $a$ to $b$ and columns $c$ to $d$, and $\mathbf{X}\succeq 0$ implies that $\mathbf{X}$ is positive semi-definite. $\mathbf{I}$ and $\mathbf{0}$ denote an identity matrix and an all-zero matrix, respectively, with appropriate dimensions. The distribution of a circularly symmetric complex Gaussian (CSCG) random vector with zero mean and covariance matrix $\bm{\Sigma}$ is denoted by $\mathcal{N}_{c}(\bm{0},\bm{\Sigma})$; and $\sim$ stands for “distributed as”.
@@ II System Model and Problem Formulation
where $\delta_{k}\triangleq\frac{\Delta}{r_{k}}$ and $n=1,2,\dots,N$. Note that $\delta_{k}\ll 1$ since the RA separation $\Delta$ is generally on the order of wavelength in practice.
@@ II-A Antenna Boresight Rotation
The initial orientations/boresights of all RAs are assumed to be parallel to the $z$-axis, and each RA’s boresight direction can be independently adjusted in 3D space by mechanical or electronic methods [43, 44, 39]. Specifically, mechanical control methods typically utilize a servo motor or micro-electromechanical system (MEMS) to physically steer the antenna boresight [2]. Note that MEMS-based actuators generally operate with milliwatt-level power consumption and millisecond-scale response times [40, 21]. In contrast, electronic control methods maintain fixed antenna orientations while enabling the antenna boresight rotation through electronic techniques such as reconfigurable parasitic radiator element loading or PIN diode switching, achieving much faster response time at the microsecond scale [42, 31, 13, 30].
@@
where $f_{x,n}$, $f_{y,n}$, and $f_{z,n}$ are the projections of RA $n$’s pointing vector on the $x$-, $y$-, and $z$-axes, respectively. For the boresight rotation of RA $n$, we let $\theta_{\mathrm{z},n}$ represent its zenith angle (i.e., the angle between the boresight direction of RA $n$ and the $z$-axis) and $\theta_{\mathrm{a},n}$ represent its azimuth angle (i.e., the angle between the projection of the boresight direction of RA $n$ onto the $x$-$y$ plane and the $x$-axis). Accordingly, we have $f_{x,n}=\operatorname{sin}\theta_{\mathrm{z},n}\operatorname{cos}\theta_{\mathrm{a},n}$, $f_{y,n}=\operatorname{sin}\theta_{\mathrm{z},n}\operatorname{sin}\theta_{\mathrm{a},n}$ and $f_{z,n}=\operatorname{cos}\theta_{\mathrm{z},n}$. Furthermore, we have $\|\vec{\mathbf{f}}_{n}\|=1$ due to normalization. To account for practical rotational constraint and mitigate antenna coupling between any two RAs, the zenith angle of each RA is confined to a specific range [11]:
@@
where $\theta_{\mathrm{max}}\in[0,\frac{\pi}{2}]$ is the maximum zenith angle that each RA is allowed to adjust.
@@ II-B Channel Model
We consider the scattering environment with $Q$ distributed scatterer clusters, where the position of scatterer cluster $q$ is represented by $\mathbf{c}_{q}\in\mathbb{R}^{3\times 1}$ with $q=1,2,\dots,Q$. In the following, we consider the narrow-band frequency-flat channel model for ease of exposition. The channel model will be extended to incorporate polarization effects and frequency-selective fading in Appendix E, with corresponding simulation results presented in Section V-C.
Based on the Friis Transmission Equation and the directional gain pattern adopted in (5), the channel power gain between user $k$ and RA $n$ can be modeled as [9]
$g_{\mathrm{U},k}(\vec{\mathbf{f}}_{n})$ $\approx\int_{\mathbb{A}_{n}}\frac{1}{4\pi\|\mathbf{u}_{k}-\mathbf{a}\|^{2}}G_{0}\left[\frac{\vec{\mathbf{f}}_{n}^{T}(\mathbf{u}_{k}-\mathbf{a})}{\|\mathbf{u}_{k}-\mathbf{a}\|}\right]_{+}^{2p}\mathrm{d}\mathbf{a}$
$=\frac{A}{4\pi(r_{k,n}^{\mathrm{B-U}})^{2}}G_{0}\left[\operatorname{cos}(\epsilon_{k,n})\right]_{+}^{2p},$ (6)
where $x]_{+}\triangleq\operatorname{max}\{x,0\}$ denotes the positive operator, the integral space $\mathbb{A}_{n}=\left[n_{x}\Delta-\frac{\sqrt{A}}{2},n_{x}\Delta+\frac{\sqrt{A}}{2}\right]\times\left[n_{y}\Delta-\frac{\sqrt{A}}{2},n_{y}\Delta+\frac{\sqrt{A}}{2}\right]$ corresponds to the surface region of RA $n$, $\mathbf{a}\in\mathbb{A}_{n}$ represents any point on plane $\mathbb{A}_{n}$, and $\operatorname{cos}(\epsilon_{k,n})\triangleq\vec{\mathbf{f}}_{n}^{T}\vec{\mathbf{u}}_{k,n}$ is the projection between $\vec{\mathbf{f}}_{n}$ and $\vec{\mathbf{u}}_{k,n}$ with $\vec{\mathbf{u}}_{k,n}\triangleq\frac{\mathbf{u}_{k}-\mathbf{w}_{n}}{\|\mathbf{u}_{k}-\mathbf{w}_{n}\|}$ being the direction vector from RA $n$ to user $k$. Similarly, the channel power gain between scatterer cluster $q$ and RA $n$ is modeled as
$g_{\mathrm{C},q}(\vec{\mathbf{f}}_{n})\approx\frac{A}{4\pi(r_{q,n}^{\mathrm{B-C}})^{2}}G_{0}\left[\operatorname{cos}(\tilde{\epsilon}_{q,n})\right]_{+}^{2p},$ (7)
where $r_{q,n}^{\mathrm{B-C}}=\|\mathbf{c}_{q}-\mathbf{w}_{n}\|$ is the distance between scatterer cluster $q$ and RA $n$, and $\operatorname{cos}(\tilde{\epsilon}_{q,n})\triangleq\vec{\mathbf{f}}_{n}^{T}\vec{\mathbf{c}}_{q,n}$ is the projection between $\vec{\mathbf{f}}_{n}$ and $\vec{\mathbf{c}}_{q,n}$ with $\vec{\mathbf{c}}_{q,n}\triangleq\frac{\mathbf{c}_{q}-\mathbf{w}_{n}}{\|\mathbf{c}_{q}-\mathbf{w}_{n}\|}$ denoting the direction vector from RA $n$ to scatterer cluster $q$. Note that the channel power gains modeled in (6) and (7) account for both path gain and directional gain, while the effect of polarization is incorporated in (69) and (70) of Appendix E.
@@
$h_{k}^{\mathrm{NLoS}}(\vec{\mathbf{f}}_{n})$ $=\sum_{q=1}^{Q}{\frac{\sqrt{\varsigma_{q}{g_{\mathrm{C},q}(\vec{\mathbf{f}}_{n})}}}{r_{k,q}^{\mathrm{U-C}}}}e^{-j\frac{2\pi}{\lambda}(d_{q,n}+o_{k,q})+j\chi_{q}},$ (9)
where $\lambda$ is the signal wavelength, $\varsigma_{q}$ represents the radar cross section (RCS) of scatterer cluster $q$, $\chi_{q}$ represents the phase shift introduced by scatterer cluster $q$, and $r_{k,q}^{\mathrm{U-C}}=\|\mathbf{u}_{k}-\mathbf{c}_{q}\|$ denotes the distance between user $k$ and scatterer cluster $q$. Thus, by superimposing the LoS and NLoS channel components, the overall multipath channel between user $k$ and the BS is given by [^1] The current geometric channel model can be refined by incorporating the standardized path loss models defined in 3GPP TR 38.901 [1] and extended to account for frequency-selective fading, as shown in Appendix E.